OpenStudy (anonymous):

3^X+3 - 9^2X-1=0 what are the steps to finding the solution?

myininaya (myininaya):

\[3^x+3-9^{2x}-1=0?\]

myininaya (myininaya):

Or! Do you mean \[3^{x+3}-9^{2x-1}=0\]

OpenStudy (anonymous):

second one

myininaya (myininaya):

Ok did you add to both sides \[9^{2x+1}\]

myininaya (myininaya):

\[3^{x+2}=9^{2x+1}\]

myininaya (myininaya):

oops - not + oh well

myininaya (myininaya):

Do you know how to answer the following What number satisfies the equation \[9=3^{?}\]

myininaya (myininaya):

Ok great. So we have \[3^{x+2}=(3^{2})^{(2x-1)}\] since 3^2=9 :)

myininaya (myininaya):

So we have the same base on both sides All you need to do is set the exponents equal and solve for x

myininaya (myininaya):

\[x+2=2(2x-1)\]

myininaya (myininaya):

Let me know if you have further questions :)

OpenStudy (anonymous):

firstly its x+3 but where did you get the 2 in "2"(2x-1)

myininaya (myininaya):

\[9=3^2 \] \[3^{x+3}=9^{2x-1}\] \[3^{x+2}=(3^2)^{(2x-1)}\]

OpenStudy (anonymous):

how does 3^2(^2x-1) = 2(2x-1)

myininaya (myininaya):

No....we have x+2=2(2x-1) not what you wrote